OpenStudy (anonymous):

Rachel's average score after 6 tests is 83. If rachel earns a score of 97 on the 7th test, what is her new average?

OpenStudy (anonymous):

\[83=\frac{ x }{ 6 }....>x=83\times6\] \[The..new,,average=\frac{ 97+83\times6 }{ 7 }=85\]